OpenStudy (anonymous):

What is fourth proportional to 3,9,21? A- 63 B- 27 C- 81 D- 54

OpenStudy (♪chibiterasu):

This is a geometric sequence. We can divide one of the values by the one before it. i.e. 21 and 9. 21/9 = 3 Each term is increasing by a factor of 3. 3*3 = 9 9*3 = 21 So what is 21*3?

OpenStudy (anonymous):

9*3=27

OpenStudy (♪chibiterasu):

Hm.

OpenStudy (♪chibiterasu):

Does the question look like this? 3;9 = 21;x

OpenStudy (♪chibiterasu):

Or just like 3;9;21;x

OpenStudy (anonymous):

3,9,21,? like this

OpenStudy (♪chibiterasu):

Nevermind. We just "cross-multiply" 3;9;21;x 3x = 21(9) 3x = 189 3x/3 = 189/3 x = ?

OpenStudy (♪chibiterasu):

Yeah I thought this was a geometric sequence, I was wrong. Hehe. ^That should be the right way to do it even though I came out with the same thing.

OpenStudy (anonymous):

i think 3*2=6 6+3=9 second is 9*2=18 18+3=21 tried is not their options 21*2=42 42+3=45

OpenStudy (anonymous):

45 is not right

OpenStudy (♪chibiterasu):

Where did you get all that from?

OpenStudy (anonymous):

in ps exam books?

OpenStudy (usukidoll):

I know this doesn't show up in higher level math , so that's a good sign

OpenStudy (anonymous):

which one is right option